Frozen15:

What value of y makes the equation true?    -7/12 = y/36 A. -28 B. -21 C. 21 D. 28

Frozen15:

@HuskyNation

alones:

Wouldn't the answer be \(y=−21\)

Frozen15:

Hmmm i was thinkin that but wasnt totally sure

alones:

Since you cross multiply first \[\frac{ -7 }{ 12 }=\frac{ y }{ 36}\] \((−7)*(36)=y*(12)\) \(−252=12y\)

alones:

Than \(flip\) the equation: \(12y=−252\) and divide both sides by \(12.\) \[\frac{ 12y }{ 12 }=\frac{ −252 }{ 12 }\] Which would end up as \(-21\) (:

Frozen15:

Ohh gotcha. thanks! Can you help me some more?

alones:

lol sure xD

HuskyNation:

sry wifi was down

Frozen15:

your fine :)

Frozen15:

What value of q makes the equation true? 18/q = 3/5 A. 6 B. 18 C. 25 D. 30

Frozen15:

@alones @HuskyNation

HuskyNation:

D

Frozen15:

thanks!

Frozen15:

Lupe played six tennis matches last weekend and won four of them.    What fraction of the matches did Lupe win?   A. 1/3   B. 1/2   C. 2/3   D. 3/4

HuskyNation:

C

Frozen15:

ty

Frozen15:

There are 16 girls and 12 boys on the cross country team.    What fraction of the team is made up of boys?   A. 3/7   B. 4/7   C. 2/3   D. 3/4 Last question

Justin234:

3*4/7*4=3/7 A. 3/7

Frozen15:

Ahhh makes sense thank you :P

Justin234:

yup np :D

Frozen15:

:)

Frozen15:

that was correct huzza!!

Justin234:

yay good for u